# On multiplicities in length spectra of semi-arithmetic hyperbolic surfaces

### v1

@@ 3. Results
When restricted to $\Omega$, the map $G$ is $\sqrt{1+(1-\delta)^{2}}$-Lipschitz for some $10<\delta<$ 1 according to the Schwarz–Pick lemma (cf. the proof of Lemma 3.1).
Now we can apply Lemma 3.5 to find a $\sqrt{1+(1-\delta)^{2}}$-Lipschitz map $G_{1}:S\to X$. Its equivariant lift $\tilde{G}:\mathbb{H}\to\tilde{X}\subset\pm\mathbb{H}\times\pm\mathbb{H}$ is also Lipschitz with the same Lipschitz constant as $G_{1}$.
Repeating the argument of Lemma 3.1, we consider $\tilde{\Omega}$, a connected lift of $\Omega$ to $\mathbb{H}$, and for any hyperbolic element $\gamma\in\Gamma,$ up to conjugation, we can assume that the displacement $\ell$ of $\gamma$ is realized by a point $z\in\tilde{\Omega},$ which means that $\ell=d(z,\gamma z).$ We know that $\rho(\gamma)$ is a hyperbolic element of $\tilde{\Gamma}$, with displacement $\sqrt{\ell^{2}+\ell_{\sigma}^{2}},$ where $\ell_{\sigma}$ is the displacement of $\gamma^{\sigma}.$ By the choice of $z$ and the fact that $\tilde{G}(w)=(w,F(w))$ for all $w\in\tilde{\Omega}$, we obtain
$\ell^{2}+\ell_{\sigma}^{2}\leq d\Big{(}(z,F(z)),\rho(\gamma)(z,F(z))\Big{)}^{2}=d(z,\gamma z)^{2}+d(F(z),\gamma^{\sigma}F(z))^{2}.$
Moreover,
$d\Big{(}(z,F(z)),\rho(\gamma)(z,F(z))\Big{)}^{2}=d(\tilde{G}(z),\tilde{G}(\gamma z))^{2}\leq\big{(}1+(1-\delta)^{2}\big{)}d(z,\gamma z)^{2},$
which implies, using the equality $\ell=d(z,\gamma z)$, that
(9) $\ell_{\sigma}\leq d(F(z),\gamma^{\sigma}F(z))\leq(1-\delta)\ell.$
@@
There are two possibilities, the first one occurs when $\ell_{\sigma}=0$, i.e. $\gamma^{\sigma}$ is elliptic, in which case the lemma follows immediately. If $\gamma^{\sigma}$ is hyperbolic, then we have $\ell_{\sigma}=2\operatorname{arcosh}\left(\frac{|\sigma(\mathrm{tr}(\gamma))|}{2}\right)$. By (9), we obtain
@@
$2\operatorname{arcosh}\left(\frac{|\sigma(\mathrm{tr}(\gamma))|}{2}\right)\leq 2(1-\delta)\operatorname{arcosh}\left(\frac{|\mathrm{tr}(\gamma)|}{2}\right),$
and the proof finishes as in Lemma 3.1.
@@
It remains to consider the case when the group $\Gamma$ has torsion. Then we first prove the inequality for its finite index torsion-free subgroup $\Gamma_{1}$, $[\Gamma:\Gamma_{1}]=q$, derived from a quaternion algebra. By formula (1), the inequality extends to $\Gamma$ with the new multiplicative constant $C_{1}=C_{1}(\Gamma_{1},q)>0$. ∎
@@
Let $\Gamma$ be a non-cocompact Fuchsian group satisfying the assumptions. By Lemma 3.6, the group $\Gamma$ satisfies the same weak bounded clustering property of Lemma 3.2. The rest of the argument follows the line of the proof of Theorem 3.4. ∎
@@ 4.1. Triangle group examples
Consider a geodesic triangle in $\mathbb{H}$ with angles $\pi/a,\pi/b,\pi/c$, where $a,b,c$ are positive integers (possibly infinite). Such a triangle always exists and is unique up to isometry as long as $11/a+1/b+1/c<$ 1. Reflections across the geodesic lines supporting the sides of the triangle generate a discrete subgroup of the full group of isometries of $\mathbb{H}$. The index $2$ subgroup of orientation-preserving isometries is a Fuchsian group of finite coarea and presentation $\langle x,y,z\mid x^{a}=y^{b}=z^{c}=xyz\rangle$, called a triangle group.
@@ 4.2. Veech group examples
Let $N$ be the subgroup of $\mathrm{Stab}(\Delta_{q})$ consisting of those elements that fix $\Delta_{q}$ pointwise. In particular, elements of $N$ are (conformal) automorphisms of $X$, and so $N$ is finite. Note that each $h\in\mathrm{Stab}(\Delta_{q})$ corresponds to an isometric automorphism of $\mathbb{H}$, and thus to an element of $\mathrm{PSL}(2,\mathbb{R})$. We may then realise the quotient $N\backslash\mathrm{Stab}(\Delta_{q})$ as a subgroup of $\mathrm{PSL}(2,\mathbb{R})$, denoted $\mathrm{PSL}(X,q)$.
Alternatively, define $\mathrm{Aff}^{+}(q)$ to be the group of quasi-conformal self/homeomorphisms $g$ of $X$ that are affine with respect to the flat structure induced by $q$. This means that, for each $p\in X\setminus\{\text{zeros of }q\}$ there are preferred coordinate charts $\zeta_{1}$ around $p$ and $\zeta_{2}$ around $g(p)$ such that $\zeta_{2}\circ g\circ\zeta_{1}^{-1}$ is an affine homeomorphism between domains of $\mathbb{R}^{2}$. The derivative of this affine homeomorphism is independent of $p$ and of the coordinate charts up to sign. Its determinant has norm 1, since $\lVert q\rVert=1$. This induces a homomorphism $a:\mathrm{Aff}^{+}(q)\to\mathrm{PSL}(2,\mathbb{R})$, taking an affine quasi-conformal homeomorphism $g$ to its derivative $a(g)$ with respect to the flat structure of $q$.
@@ 5. Quantum chaos
Consider the motion of a free point particle on an oriented surface of constant negative curvature $\𝑆\ΓS=\Gamma\backslash\mathbb{H}$. This is a classical Hamiltonian system whose trajectories are defined by the geodesic flow on the cotangent bundle $T^{*}(S)$. It is well known that the geodesic flow is chaotic, meaning that when restricted to a fixed energy level (a time-independent Hamiltonian) the flow is ergodic and has almost everywhere positive Lyapunov exponents [AA68]. Its trajectories are very sensitive to the initial data and unpredictable in this sense. At the opposite extreme we have the integrable systems whose time evolution in the phase space $T^{*}(S)$ takes place on invariant tori. These systems are the most regular and predictable.
@@
$-\hslash^{2}\Delta\psi_{j}=\lambda_{j}\psi_{j}.$
@@
[…]

### v2

@@ 3. Results
When restricted to $\Omega$, the map $G$ is $\sqrt{1+(1-\delta)^{2}}$-Lipschitz for some $0<\delta<1$ according to the Schwarz–Pick lemma (cf. the proof of Lemma 3.1).
By applying Lemma 3.5, we can find a $\sqrt{1+(1-\delta)^{2}}$-Lipschitz map $G_{1}:S\to X$ which is homotopic to $G$.
Given any hyperbolic element $\gamma\in\Gamma,$ we consider the closed geodesic $\overline{\gamma}\subset S$ of length $\ell$ given by the projection of the axis of $\gamma$. By the definition of $G$, the image $G(\gamma)$ is freely homotopic to the projection on $X$ of the axis of the hyperbolic element $\rho(\gamma)\in\Delta$, whose length is given by $\sqrt{\ell^{2}+\ell_{\sigma}^{2}}$, where $\ell_{\sigma}$ is the displacement of $\gamma^{\sigma}.$ Since $G$ is homotopic to $G_{1}$, we conclude that $G(\gamma)$ is homotopic to $G_{1}(\gamma)$. Hence, the displacement of $\rho(\gamma)$ is smaller than the length of the closed curve $G_{1}(\gamma)$. Now we use the fact that $G_{1}$ is $\sqrt{1+(1-\delta)^{2}}$-Lipschitz in order to get the estimate
$\ell^{2}+\ell_{\sigma}^{2}\leq\big(1+(1-\delta)^{2}\big)\ell^{2}.$
Hence,
(9) $\ell_{\sigma}\leq(1-\delta)\ell.$
@@
There are two possibilities, the first one occurs when $\ell_{\sigma}=0$, i.e. $\gamma^{\sigma}$ is elliptic, in which case the lemma follows immediately. If not, then $\gamma^{\sigma}$ is hyperbolic and we have $\ell_{\sigma}=2\operatorname{arcosh}\left(\frac{|\sigma(\mathrm{tr}(\gamma))|}{2}\right)$. By (9), we obtain
@@
$2\operatorname{arcosh}\left(\frac{|\sigma(\mathrm{tr}(\gamma))|}{2}\right)\leq 2(1-\delta)\operatorname{arcosh}\left(\frac{|\mathrm{tr}(\gamma)|}{2}\right);$
$|\sigma(\mathrm{tr}(\gamma))|\leq 2\cosh\left((1-\delta)\operatorname{arcosh}\left(\frac{|\mathrm{tr}(\gamma)|}{2}\right)\right).$
This is estimate (6) from Lemma 3.1 and the proof concludes in the same way.
@@
2. It remains to consider the case when the group $\Gamma$ has torsion. Let us fix a finite index torsion-free subgroup of the group $\Gamma^{(2)}$ and denote it $\Gamma_{1}$.
With the group $\Gamma_{1}$ chosen, by (9) we have that $\ell(\beta^{\sigma})\leq(1-\delta)\ell(\beta)$ for some constant $\delta>0$ and all hyperbolic elements $\beta\in\Gamma_{1}$. Now we notice that this inequality can be extended to all hyperbolic $\gamma\in\Gamma^{(2)}$. Indeed, given any such $\gamma$, there exists a positive integer $j$ such that $\gamma^{j}\in\Gamma_{1}$. Since the action of $\sigma$ preserves the products, we have $(\gamma^{\sigma})^{j}=(\gamma^{j})^{\sigma}$. By the previous argument, the length of the geodesic associated to $(\gamma^{j})^{\sigma}$ satisfies inequality (9). Hence for any hyperbolic $\gamma\in\Gamma^{(2)}$, we have
$\ell(\gamma^{\sigma})=\frac{\ell((\gamma^{j})^{\sigma})}{j}\leq\frac{(1-\delta)\ell(\gamma^{j})}{j}=(1-\delta)\ell(\gamma).$
This is an extension of formula (9) to the group $\Gamma^{(2)}$ and the proof concludes as in part 1. ∎
@@
Let $\Gamma$ be a non-cocompact Fuchsian group satisfying the assumptions. By Lemma 3.6, the group $\Gamma$ satisfies the weak bounded clustering property of Lemma 3.2. The rest of the argument follows the line of the proof of Theorem 3.4. ∎
@@ 4.1. Triangle group examples
Consider a geodesic triangle in $\mathbb{H}$ with angles $\pi/a,\pi/b,\pi/c$, where $a,b,c$ are positive integers (possibly infinite). Such a triangle always exists and is unique up to isometry as long as $1/a+1/b+1/c<1$. Reflections across the geodesic lines supporting the sides of the triangle generate a discrete subgroup of the full group of isometries of $\mathbb{H}$. The index $2$ subgroup of orientation-preserving isometries is a Fuchsian group of finite coarea and presentation $\langle x,y,z\mid x^{a}=y^{b}=z^{c}=xyz\rangle$, called a triangle group.
@@ 4.2. Veech group examples
[…]
Alternatively, define $\mathrm{Aff}^{+}(q)$ to be the group of quasi-conformal self-homeomorphisms $g$ of $X$ that are affine with respect to the flat structure induced by $q$. This means that, for each $p\in X\setminus\{\text{zeros of }q\}$ there are preferred coordinate charts $\zeta_{1}$ around $p$ and $\zeta_{2}$ around $g(p)$ such that $\zeta_{2}\circ g\circ\zeta_{1}^{-1}$ is an affine homeomorphism between domains of $\mathbb{R}^{2}$. The derivative of this affine homeomorphism is independent of $p$ and of the coordinate charts up to sign. Its determinant has norm 1, since $\lVert q\rVert=1$. This induces a homomorphism $a:\mathrm{Aff}^{+}(q)\to\mathrm{PSL}(2,\mathbb{R})$, taking an affine quasi-conformal homeomorphism $g$ to its derivative $a(g)$ with respect to the flat structure of $q$.
@@ 5. Quantum chaos
[…]
@@
$\hslash^{2}\Delta\psi_{j}=\lambda_{j}\psi_{j}.$
@@
$\frac{1}{L}\int_{0}^{L}\Sigma^{2}(\lambda,\xi)d\xi\gg\frac{\sqrt{\lambda}}{\log^{2}(\lambda)},\text{ for }\frac{\sqrt{\lambda}}{\log(\lambda)}\ll L\ll\sqrt{\lambda}.$
@@
Data Availability. Data sharing is not applicable to this article as no datasets were generated or analysed during the current study.
Declarations
Conflict of interest. The authors have no Conflict of interest to declare that are relevant to the content of this article.